OpenStudy (anonymous):

When the following quadratic equation is written in general form, what is the value of "c"? -3/4x^2 + 2 = 0 is it -2?

OpenStudy (michele_laino):

the general form of a quadratic equation is: \[a{x^2} + bxy + c{y^2} + dx + ey + f = 0\]

OpenStudy (michele_laino):

where a, b, c, d, e, and f are real parameters

OpenStudy (anonymous):

woah. your talking about quadratics?

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

if we consider a quadratic equation in one variable: \[a{x^2} + bx + c = 0\] then we can rewrite your equation as follows: \[\frac{3}{4}{x^2} - 2 = 0\]

OpenStudy (michele_laino):

or: \[3{x^2} - 8 = 0\]

OpenStudy (anonymous):

add 8 divide by 3? or no

OpenStudy (michele_laino):

no, since we have rewritten your equation using integers parameters

OpenStudy (michele_laino):

so, I think that a possible choice, can be this: a=3, b=0, c=-8

OpenStudy (anonymous):

thanks i got the answer c = -8 :)

OpenStudy (michele_laino):

:)